### v1

# references.bib

Abstract
As the burden of herbicide resistance grows and the environmental repercussions of excessive herbicide use become clear, new ways of managing weed populations are needed. This is particularly true for cereal crops, like wheat and barley, that are staple food crops and occupy a globally significant portion of agricultural land. Even small improvements in weed management practices across these major food crops worldwide would yield considerable benefits for both the environment and global food security. Blackgrass is a major grass weed which causes particular problems in cereal crops in north-west Europe, a major cereal production area, because it has high levels of of herbicide resistance and is well adapted to agronomic practice in this region. With the use of machine vision and multispectral imaging, we investigate the effectiveness of state-of-the-art methods to identify blackgrass in wheat and barley crops. As part of this work, we provide a large dataset with which we evaluate several key aspects of blackgrass weed recognition. Firstly, we determine the performance of different CNN and transformer-based architectures on images from unseen fields. Secondly, we demonstrate the role that different spectral bands have on the performance of weed classification. Lastly, we evaluate the role of dataset size in classification performance for each of the models trialled. We find that even with a fairly modest quantity of training data an accuracy of almost 90% can be achieved on images from unseen fields.
[…]
1 Introduction
Weeds compete with crops for light, water and nutrients. Additionally, grains from grass weeds get harvested alongside grains from the crop, contaminating the harvest, and can drastically lower the price received for the crop. Current mainstream weed management practices involve broadcast spraying, where the entire field is sprayed with a selective herbicide. An alternative approach, known as precision spraying, targets herbicide only at areas of weed cover guided by machine vision. Precision spraying facilitates up to 95% reduction in herbicide use (depending on crop and field) [zanin2022reduction], and a proportional decrease in herbicide costs and environmental damage from runoff. In addition, precision spraying allows a wider range of herbicides (potentially broad spectrum) to be used, opening more options for resistance management [busi2020rotations]. Meanwhile, other technologies, like automated camera-guided mechanical weeding, promise non-chemical weed control. Both of these promising technologies are predicated on accurate, real-time weed recognition within often complex crop canopies.
Efforts to reduce herbicide usage in staple cereal crops have the potential to deliver significant impact. The area of global cropland devoted to growing cereals, and in particular rice, wheat and maize, is orders of magnitude greater than that of many vegetable crops [fao]. Furthermore, wheat is second only to rice as a global staple [Eren2022], with a global consumption of 65.6 kg per person per year. Therefore, not only is sustainable and reliable crop care for wheat critical for global food security, any technology that reduces herbicide usage in these global staples has the potential to massively reduce the environmental impact of herbicide use globally.
Moreover, particular weeds present a greater threat to global food supply than others; grass weeds are a particular problem in wheat production due to their biological similarities [chhokar2012weed, young1996weed, moyer1994weed, Hick2018]. In Europe, blackgrass is one of the most economically damaging weeds [Vara2020], so effective strategies to manage populations are a priority. Due to high levels of resistance to selective herbicides [Hick2018], using them for precision spraying may not be sufficient to control blackgrass populations. In these instances, precision spraying with non-selective herbicides could be an alternative approach, but requires a high degree of confidence in detection and targeting so as not to damage the crop. Such confidence will rely on well tested, accurate, real time machine vision systems that can be deployed at scale to the field.
2 Background
We start by reviewing the relevant research. First we look at how the variety of color spaces have been used in the research to date, and secondly, how those color spaces, along with other features, have been used to recognise weeds. Lastly, we systematically evaluate all recent research on weed recognition in wheat and barley to understand on which weed species research effort has been focused.
2.1 Color Spaces for Weed Recognition
Initially, color indices derived from RGB like Excess Greenness (ExG), Excess Redness (ExR), ExG minus ExR (ExGR) and Normalized Difference Index (NDI) were used in weed recognition research [jensen2021predictive]. Multispectral imaging for crop and weed discrimination is now an active area of research. There is an existing precedent for the use of multispectral imaging in plant analysis since the Normalized Difference Vegetation Index (NDVI) [rouse1974monitoring], derived from near-infrafed (NIR) and red light, has long been used in agriculture to assess crop health. Studies trialling different combinations of NIR, Red and NDVI as inputs to a convolutional neural network (CNN) to detect weeds in sugar beet fields found that NIR and Red values used together performed the best when identifying weeds [sa2018weednet, razaak2019integrated]. Interestingly, including NDVI as an input, alongside NIR and Red values, did not improve the performance [sa2018weednet, razaak2019integrated]. With the use of hyperspectral cameras, it was found that increasing the number of bands from 3 up to 61 resulted in incremental improvements in weed species classification accuracy [farooq2018analysis]. In a study where several color indices where trialled, the inclusion of NIR information either alone, or alongside other indices like ExG and NDI, significantly improved the segmentation accuracy [wang2020semantic].
2.2 Weed Recognition Techniques
Early research into techniques for weed recognition relied on hand-crafted feature-based techniques. Approaches to discriminating grass weeds in cereal crops use color [anderegg2023farm], texture [xu2020recognition], depth [xu2020recognition] and spectral [amziane2021weed, jensen2021predictive] features. For weed classification, support vector machine (SVM) [cristianini2000introduction] and k-Nearest Neighbor (k-NN) [cover1967nearest] approaches were used to learn decision boundaries in the feature space [xu2020recognition].
Deep learning-based weed classification rose to prominence after the success of early CNN architectures like AlexNet [krizhevsky2012imagenet] and VGG [simonyan2014very]. However, ResNet [he2016deep] improved the applicability of this technology by utilising skip connections to address the vanishing gradient problem, which enabled the training of deeper neural networks, resulting in improved performance and convergence. In the following years, deep learning-based methods with faster inference speeds and smaller memory footprints enabled weed recognition on the edge, like the popular MobileNetV3 architecture, which exploits lightweight depthwise separable convolutions for improved efficiency [howard2019searching]. EfficientNet employed a neural architecture search approach to design a network and a novel method to scale the model depth, width, and resolution of the network in order to balance accuracy and efficiency[tan2019efficientnet]. Recently, transformers have demonstrated how powerful self-attention mechanisms can be for recognizing complex patterns in sequential data [vaswani2017attention]. Vision Transformer (ViT) [dosovitskiy2020image] applied transformer technology to image classification by treating an image as a sequence of patches, which were then linearly embedded and processed through a transformer architecture. The self-attention mechanism in transformers allowed ViT to analyze global contextual information, improving performance on various image classification benchmarks.
ResNet, MobileNetV3 and MobileViT [mehta2021mobilevit] — a compact version of ViT — demonstrated strong performance in weed classification in wheat [alirezazadeh2023comparative]. Deep learning-based object detection techniques, that identify the region where an object is located in an image, have been used to detect grass weeds in wheat crops including using DetectNet [dyrmann2017roboweedsupport], SSD [dyrmann2018using], Faster R-CNN [xu2021multi, xu2023weedsnet]. Additionally, there is work segmenting grass species, such as ryegrass, using an adapted ERFNet [su2021real], as well as green foxtail and horsetail using an adapted U-Net network [zou2021modified].
2.3 Prevalence of Grass Weed Recognition in Cereal Crops
We systematically reviewed the latest machine vision research on wheat and barley crops and identified that grass weeds are understudied compared to broadleaf weeds. Grass weeds are more challenging to distinguish from cereal crops (which are themselves grass species). Papers were identified in Google Scholar and Scopus using search query “weed identification” OR “weed recognition” OR “weed discrimination” OR “weed detection” OR “weed classification” OR “weed segmentation” on 20th October 2023. All publications from recognised, peer-reviewed, English-language journals and conference venues that could be accessed were included. For this analysis, we are only interested in weed detection in wheat and barley crops, so excluding all other weed detection papers we found a total of 79 papers.
Our analysis of the distribution of weeds species used in weed recognition papers in wheat and barley crops is shown in Figure 1. Each weed species from each paper is added to the cumulative total for that species. The weed species are divided into eudicots (broadleaf plants) and monocots (grasses). While the common names are used in the figure, the full list of papers categorised by weeds, along with their Latin names, is available in Appendix A.
Figure 1 shows that broadleaf weeds have garnered more research interest than grass weeds. One possible explanation for this discrepancy is how challenging weed recognition is in grass weeds compared to broadleaf weeds. Grass crops also present a high level of challenge compared to broad leaf crops as they occlude one another, as well as any weeds in the field, even at relatively early growth stages. By contrast, many broadleaf crops remain spaced out, display distinctive features, for example distinctive color and leaf shape, at early growth stages. Additionally, many grass crops look very similar to other grass plants. Distinguishing between grass species is an example of fine-grained visual classification [ma2021research]. This is a subgroup of classification tasks that aims to distinguish between similar examples based on some subcategory like plant species or car model. Moreover, like with the classification of medical images, another fine-grained visual classification task, experts are required to produce ground-truth images — adding to the cost of producing a dataset. Therefore, grass weed recognition in cereal crops has some of the same challenges as medical imaging domains when it comes to building datasets.
Figure 1: Weed detection research effort in wheat and barley across weed Species, as measured by published papers, for broad leaf (blue) and grass (red) weeds.
(a) Wheat
(b) Blackgrass
Figure 2: Example images from the Plant Seedlings Dataset [giselsson2017public].
An early contribution to the field of deep learning for weed detection in wheat was the Plant Seedling Classification dataset [giselsson2017public]. This dataset contained images of crops including wheat, maize and sugar beet along with 9 common weed species including blackgrass. Over a third of the papers on wheat/weed recognition reviewed for this survey carried out experiments on the Plant Seedling Classification dataset [chavan2018agroavnet, binguitcha2019crops, elnemr2019convolutional, jiang2019novel, espejo2020improving, fu2020field, rahman2020performance, skacev2020development, trong2020late, espejo2021combining, espejo2021testing, farkhani2021weed, mounashree2021automated, siddiqui2021neural, trong2021yielding, damilare2022weed, kumar2022weed, kundur2022deep, mu2022faster, reddy2022weed, srivastava2022role, tara2022machine, bharathi2023plant, dheeraj2023using, guo2023identification, mishra2023comparative, mu2023densenet]. This dataset has made a great contribution to the field by drawing attention to the problem of crop and weed discrimination for crops that are critical to the world’s food supply. However, using this dataset does not enable an evaluation of in-field crop and weed discrimination.
The rest of the wheat papers use in-field images of wheat and weeds. A few papers explore the problem of weed detection in mature wheat crops using both color features [rasmussen2019pre] and LiDAR [shahbazi2021assessing]. Most of the others attempt to detect weeds at an early stage. Many of these focused solely on broadleaf weeds [pflanz2018weed, alirezazadeh2023comparative, el2022metaheuristic, zhuang2022evaluation, almalky2022efficient, almalky2023deep, almalky2023real, xu2023weedsnet] which are easier to distinguish in grass crops than grass weeds. Some of the work discriminating grasses, hand-crafted approaches using color [anderegg2023farm], texture [xu2020recognition], depth [xu2020recognition] and spectral [amziane2021weed, jensen2021predictive] features were implemented. In this work, we focus on using deep learning-based approaches as they have been shown to be more robust to changes in illumination and occlusion.
2.4 Paper Overview
To facilitate evaluating agricultural vision systems on in-context, in-field data, we present a weed image data set featuring wheat and barley of over 15,000 images across 51 fields in the UK, with and without blackgrass — a grass weed of cereal crops. This dataset is comparable in scale to the seminal Plant Seedling Classification dataset [giselsson2017public]. The visual similarity of blackgrass to cereal crops poses a challenging fine-grained classification task for precision weed management technology. The release of this dataset will accelerate the development of weed detection systems for black-grass, and due to their visual similarity, other grass weeds, for wheat, a critical global staple.
The contributions of this paper are as follows:
  • A dataset containing over 15,000 multispectral images of wheat and blackgrass stratified by field, season, geo-location and soil-type, enabling the study of domain generalisation with respect to these factors;
  • An evaluation the effect of using different spectral bands for blackgrass classification; and
  • An assessment of the effect of training data quantity on model performance providing guidelines for future dataset building.
3 Materials and Methods
In this section, we first present our dataset of blackgrass in wheat and barley crops. We then outline our detection approach including architectures, experiments, training settings and evaluation methods.
3.1 Dataset
The dataset [^1] The dataset will be made publicly available upon acceptance. consists of images of blackgrass in wheat and barley crops collected from 51 fields across eight different soil types in the east of England, a key grain growing region of the UK, between 3rd November 2020 and 16th October 2021. An average of 300 images was taken from each field. The majority of fields had more than 100 images both with and without blackgrass, however, seven fields only had ‘blackgrass’ images and four fields only had ‘no blackgrass’ images. As shown in Table 1, the images were split into by field so that there is approximately a 80-10-10 split for training, validation and test sets. The aim of this stratification is to ensure that each model’s ability to generalise to unseen fields is assessed. Table 3.1 shows how many images are in each set according class, crop type and season. Care was taken to ensure the sets were balanced in terms of positive and negative examples as well as to ensure each set contain some images from each stage in the season and each crop type.
Dataset Number of Fields
Train 37
Validation 8
Test 6
Table 1: Training, validation and testing is performed on separate sets of images from different fields.
{tblr}
—c—c—c—c—c—c— Train Validation Test Total
[…]
dashed]2-6 No Blackgrass 5982 1036 1023 8041
[…]
dashed]2-6 Spring Barley 542 123 372 1037
dashed]2-6 Winter Barley 1613 205 177 1995
[…]
dashed]2-6 Mid 4571 841 1166 6578
dashed]2-6 Late 3550 883 369 4802
Table 2: Stratification of image numbers for each class, crop and season.
Land managers were consulted to find patches of blackgrass in wheat and barley fields. Images of class ‘no blackgrass’ were taken near the blackgrass patches. Fields with a good number of blackgrass patches were re-visited to collect images at different growth stages of the crop and blackgrass. Blackgrass class was assessed at the time of image collection by the same observer for all images across all fields to maximise the number of images collected. The dataset therefore only has image level labels.
Figures 3(a) and 3(b) are example images from the early growing season (November to January) where the crop plants were well separated and blackgrass was present in short, singular, thin blades. Figures 3(c) and 3(d) are from February on-wards where the crop canopy exposes less bare soil and the individual blackgrass blades developed into a bushier structure.
(a) Blackgrass: Early Season
(b) No Blackgrass: Early Season
(c) Blackgrass: Late Season
(d) No Blackgrass: Late Season
Figure 3: RGB examples from the dataset. In a) blackgrass are the thin single green strands and the wheat is the wider leaf blades. Blue arrows show a few indicative examples of blackgrass, but there are many blackgrass seedlings in a)
For each field, the soil type in that field was categorised according to the Cranfield Soil and Agrifood Institute’s Soilscapes databases [soilscapes]. Due to the geography of the farms that were visited, there were no fields visited that contained predominantly peat, or predominantly chalky soil. Table 3 describes each soil type in the dataset and shows the break down of images with each soil type in each of the dataset splits. Note that, due to prioritising a stratification based on geo-location, it was not possible to have every soil type represented in each dataset split.
Soil Type Description Train Validation Test Total
Type 1 Free draining lime-rich loamy 379 0 0 379
Type 2 Free draining, base-rich, slightly acidic loamy 497 0 581 1078
Type 3 Wet slightly acidic loamy and clayey 489 0 492 981
Type 4 Wet lime-rich loamy and clayey 1724 0 0 1724
Type 5 Acidic sandy and loamy 0 0 0 0
Type 6 Base-rich loamy and clayey, wet 1444 214 0 1658
Type 7 Loamy and clayey wet 6204 1101 197 7502
Type 8 Wet loamy and sandy, peaty surface 1015 744 848 2607
Table 3: Soil types
3.2 Neural Network Architectures
For this study, we choose ResNet-50, EfficientNet B4 and Swin-B as the representative architectures for the benchmarks. They are detailed in Table 4. ResNet [he2016deep] introduced residual learning by incorporating skip connections, enabling the training of very deep neural networks. Despite being the oldest of the architectures, it is still widely used, due to its training efficiency and accuracy. EfficientNet [tan2019efficientnet] improved image classification on ImageNet by optimizing for network size and computational efficiency using a neural architecture search approach. This was selected as another popular choice of CNN architecture but more explicitly designed for computational efficiency than ResNet. Given, the recent interest in transformers for machine vision, we chose the Swin Transformer [liu2021swin] which improved on Vision Transformer (ViT) by introducing a hierarchical architecture that processes images at multiple scales in an efficient manner. This hierarchical design facilitated capturing both local and global features effectively, addressing the limitations of the single-level processing in ViT.
Model Params GFLOPS ImageNet Acc@1 ImageNet Acc@5
Resnet 50 25.6M 4.09 76.13 92.86
Efficient Net B4 19.3M 4.39 83.38 96.59
Swin B 87.8M 15.43 83.58 96.64
Table 4: Models
3.3 Experiments
Using these models, we perform the following experiments on our dataset:
  • Baseline Models We establish baseline results for our dataset on our selected models. Furthermore, we examine the performance on each of the separate fields in the test to understand the strengths and weaknesses of the models in different environments.
  • Spectral Band Importance To gain insight into the role of each spectral band, the models were trained on all the combinations of spectral bands in the dataset.
  • Training Data Quantity Experiment To assess the significance of the training data quantity for each tested model, we evaluated them on subsets containing 25%, 50%, and 75% of the training data. These subsets were generated through random sampling from all training images across fields; maintaining the original proportion of positive and negative samples from the entire dataset.
3.4 Training Settings
The models were trained on an Nvidia GeForce RTX 2020 Ti. The initial weights for each model were taken from Pytorch’s version of the models pretrained on ImageNet. Since these models were trained on RGB input, the NIR and red edge channels were set to the pretrained weights for the red channel from these models. Each model was trained for 50 epochs. The model with the highest accuracy on the validation set is selected. The optimizer was stochastic gradient descent (SGD) [robbins1951stochastic]. The initial learning rate was 0.001 and this was decayed to 0.0001 using cosine annealing. The batch size used was 4. Cross entropy loss, also known as log loss, was used as the objective function to train the models. The training settings outlined were determined empirically.
3.5 Evaluation
In order to give a comprehensive overview of evaluation of a classification model’s performance, we use accuracy, precision, recall and MCC. Since the dataset is mostly balanced, accuracy gives a general overview of the quality of predictions. Accuracy represents that total fraction of correct predictions made by the model and is calculated as follows:
$Accuracy=\frac{TP+TN}{TP+TN+FP+FN}$ (1)
where $TP$ is the total number of true positives, $TN$ is the total true negatives, $FP$ is the total false positives and $FN$ is the total false negatives.
Additionally, precision and recall with respect to the blackgrass class offers insights into the model’s ability to correctly identify positive instances and capture all relevant positive instances, respectively. Precision is the fraction of positive predictions that are correctly predicted and is calculated as follows:
$Precision=\frac{TP}{TP+FP}$ (2)
While recall is the fraction of the positive class present in the dataset correctly predicted by the model and is calculated as follows:
$Recall=\frac{TP}{TP+FN}$ (3)
Matthews Correlation Coefficient (MCC) is another way to assess if the quality of binary classification predictions is robust to class inbalances in the dataset. While our dataset is mostly balanced, this metric allows us to account for the slight difference in class balance. Additionally, it will enable a comparison with work published on unbalanced datasets. MCC produces values between -1 and 1 where 1 indicates perfect prediction, 0 indicates a random prediction and -1 means a perfectly inverse prediction. MCC is calculated as follows:
$MCC=\frac{TP*TN-FP*FN}{\sqrt{(TP+FP)(TP+FN)(TN+FP)(TN+FN)}}$ (4)
For each of these metrics, we report the mean and 2 standard deviations of 5 training runs each with different random seeds.
4 Results
Table 5 shows the accuracy, MCC, precision and recall. Figure 4 shows the accuracy for each model on each of the different fields in the test sets. Figure 5 shows the accuracy of ResNet-50 trained on every combination of spectral band. Figure 6 shows the how the accuracy changes as the number of training images increases.
Model Accuracy MCC Precision Recall
ResNet 50 $0.873\pm 0.019$ $0.746\pm 0.019$ $0.863\pm 0.014$ $0.876\pm 0.001$
Efficient B4 $0.830\pm 0.018$ $0.661\pm 0.018$ $0.810\pm 0.014$ $0.847\pm 0.020$
Swin B $0.877\pm 0.025$ $0.756\pm 0.024$ $0.872\pm 0.029$ $0.887\pm 0.035$
Table 5: Baseline results. The mean of 5 models from repeated training runs was taken. Each metric is given with 2 standard deviations.
Table 5 shows that all models achieved an accuracy greater than 0.8. Swin B has the highest accuracy with 0.877, however, ResNet 50 performs only slightly worse with 0.873. EfficientNet B4 performs less well than the other two models with only 0.83. The models rank the same according to MCC.
Figure 4 shows that the performance of the models varied considerably between the different fields. Mostly, models performed worse on barley compared with wheat and worse on midseason crops compared to late season crops.
Figure 4: Accuracy on each field in the test set. The mean of 5 models from repeated training runs was taken. The error bars represent a range of two standard deviations around the mean.
Figure 5 highlights the importance of NIR in the discrimination of wheat and blackgrass. NIR alone outperforms RGB and RGB + NIR. However, using all spectral bands yields the highest accuracy.
Figure 5: Models trained on different combinations of color channels. Models, identified by the combination of channels used, are ordered by accuracy.
Figure 6 shows that increasing the number training images improves performance up to the first 6000 images. From there, performance does not improve. Swin-B shows more variability in performance with a small number of images than the other models.
Figure 6: Models with different quantities of training data. The mean of 5 models from repeated training runs was taken. The error bars represent a range of two standard deviations around the mean.
5 Discussion
Our study shows the need for more research into the challenging visual recognition task of detecting grass weeds in grass crops. The dataset we provide aims to offer an opportunity for further research in this area. Given the scale of the dataset, we hope it offers an opportunity to learn generalised solutions. In order to test a models ability to generalise, the validation and test set are made up of images from unseen fields — that is images from fields not in the training dataset.
Overall, the results in Table 5 show that all models achieve an average accuracy of over 80% using all the training data across all the test fields. We can conclude Swin B is the best performing on average but it is subject to more variability in performance. ResNet 50 is only slightly worse on average while EfficientNet B4 performs worse overall. The precision and recall for all models are similar, with recall only slightly higher for all models, suggesting they are quite balanced, capturing most true positives without too many false positives. By investigating the performance on each field in the test set, we find blackgrass classification accuracy improves as the crops and weeds mature in the late season.
5.1 Fields
In general, Figure 4, all models performed worse on barley crops compared with wheat crops. This could be the result of fewer images of barley crops being in the dataset. Additionally, all models performed best on the fields where images were taken during the late season. Despite fewer images in the training data of late season crops, these were more accurately classified. This suggests that as the crop and blackgrass mature they become easier to discriminate.
5.2 Channels
Moreover, the dataset is comprised of multispectral images, and allowed us to test the role different spectral bands could play in discriminating between cereal crops and blackgrass. In line with other research, we find that the addition of NIR reliably improves classification accuracy, shown in Figure 5. For example, Red on its own has an accuracy of 0.772 while Red and NIR has an accuracy of 0.860. Similarly, Red and Blue together have an accuracy of 0.770 but with NIR the accuracy is 0.884. Additionally, we find certain powerful combinations of spectral bands, Red, Blue and NIR being one such example, suggesting that particular discriminating features are present in a subset of spectral bands and that in some cases the addition of other spectral bands adds noise to the classification process. This is also shown by combinations that prove detrimental, like Green, NIR and Red Edge, that performs worse than NIR and Red Edge, and NIR alone. This indicates the features in the other colour channels contribute noise to the detection process. However, it is still the case that using all spectral bands in combination is the most accurate.
5.3 Training Data Size
Finally, since our dataset is unusually large for an agricultural classification dataset we wanted to determine whether this larger dataset conferred any benefits. We found that up to 6000 images there was an improvement in performance but beyond that there was no performance improvement, as shown in Figure 6. It could be that the variation in the 37 fields in the training set was well represented by the 6000 randomly sampled images and provided no additional discriminating information. While EfficientNet has the lowest accuracy of the approaches tried, it performs the best on a very restricted dataset size.
6 Conclusion
The problem of grass weed recognition in grass crops, including in major cereal crops, is under-represented in the research to date. To address this, we provide a large dataset of labelled images of blackgrass in wheat and barley crops and show the performance of a representative set of deep classification models in distinguishing between images with and without blackgrass. We demonstrate the performance of state-of-the-art machine vision techniques on unseen fields. We explore the usefulness of additional spectral bands in the classification as well as the importance of additional training data for different popular training models. We find NIR is important for blackgrass classification which concurs with other work that has shown it is useful for weed recognition tasks. Additionally, we establish that for our dataset and this task, there is an upper limit on the number of images that yield improvements in accuracy.
Acknowledgement
This research was partially supported by a PhD studentship from the University of Lincoln, and by Lincoln Agri-Robotics as part of the Expanding Excellence in England (E3) Programme.
Author Contributions
Madeleine Darbyshire: Conceptualization, Investigation, Methodology, Software, Visualization, Writing – original draft. Shaun Coutts: Conceptualization, Funding acquisition, Investigation, Supervision, Writing – review & editing. Eleanor Hammond: Investigation. Fazilet Gokbudak: Investigation. Cengiz Ozterli: Funding acquisition, Supervision. Petra Bosilj: Conceptualization, Funding acquisition, Methodology, Writing – review & editing. Junfeng Gao: Investigation. Elizabeth Sklar: Supervision, Writing – review & editing. Simon Parsons: Supervision, Writing – review & editing.
Appendix A Full List of Wheat and Barley Papers by Weed Species
Latin Name Common Name Total Publications
Aethusa cynapium 1 [anderegg2023farm]
Alchemilla arvensis Parsley Piert 2 [anderegg2023farm, xu2023weedsnet]
Alopecurus aequalis Shortawn Foxtail 1 [xu2022multi]
Alopecurus myosuroides Blackgrass 32 [dyrmann2015roboweedsupport, sukumar2016weed, chavan2018agroavnet, binguitcha2019crops, elnemr2019convolutional, jiang2019novel, espejo2020improving, fu2020field, rahman2020performance, skacev2020development, trong2020late, espejo2021combining, espejo2021testing, farkhani2021weed, jensen2021predictive, mounashree2021automated, siddiqui2021neural, trong2021yielding, damilare2022weed, fraccaro2022deep, kumar2022weed, kundur2022deep, mu2022faster, reddy2022weed, srivastava2022role, tara2022machine, bharathi2023plant, dheeraj2023using, guo2023identification, mishra2023comparative, mu2023densenet, rahman2023new]
Amaranthus retroflexus Redroot Pigweed 4 [xu2020recognition, ronay2022effect, xu2022multi, xu2023weedsnet]
Apera spica-venti Loose Silky-bent 30 [dyrmann2015roboweedsupport, chavan2018agroavnet, binguitcha2019crops, elnemr2019convolutional, jiang2019novel, espejo2020improving, fu2020field, rahman2020performance, skacev2020development, trong2020late, espejo2021combining, espejo2021testing, farkhani2021weed, mounashree2021automated, siddiqui2021neural, trong2021yielding, damilare2022weed, kumar2022weed, kundur2022deep, mu2022faster, reddy2022weed, srivastava2022role, tara2022machine, anderegg2023farm, bharathi2023plant, dheeraj2023using, guo2023identification, mishra2023comparative, mu2023densenet, rahman2023new]
Arctotheca calendula Cape Dandelion 1 [le2021detecting]
Avena fatua Wild Oat 1 [shahbazi2021assessing]
Bromus japonicus Japanese brome 2 [xu2022multi, xu2023weedsnet]
Capsella bursa-pastoris Shepherd’s Purse 36 [dyrmann2015roboweedsupport, dyrmann2017roboweedsupport, chavan2018agroavnet, dyrmann2018using, binguitcha2019crops, elnemr2019convolutional, jiang2019novel, espejo2020improving, fu2020field, rahman2020performance, skacev2020development, trong2020late, xu2020recognition, espejo2021combining, espejo2021testing, farkhani2021weed, mounashree2021automated, siddiqui2021neural, trong2021yielding, damilare2022weed, kumar2022weed, kundur2022deep, mu2022faster, reddy2022weed, srivastava2022role, tara2022machine, xu2022multi, zhuang2022evaluation, anderegg2023farm, bharathi2023plant, dheeraj2023using, guo2023identification, liu2023semi, mishra2023comparative, mu2023densenet, rahman2023new]
Centaurea cyanus Cornflower 1 [anderegg2023farm]
Chenopodium album Fat Hen 35 [dyrmann2015roboweedsupport, dyrmann2017roboweedsupport, chavan2018agroavnet, dyrmann2018using, wang2018low, binguitcha2019crops, elnemr2019convolutional, jiang2019novel, espejo2020improving, fu2020field, rahman2020performance, skacev2020development, trong2020late, wang2020transfer, espejo2021combining, espejo2021testing, farkhani2021weed, mounashree2021automated, siddiqui2021neural, trong2021yielding, damilare2022weed, kumar2022weed, kundur2022deep, mu2022faster, reddy2022weed, srivastava2022role, tara2022machine, alirezazadeh2023comparative, anderegg2023farm, bharathi2023plant, dheeraj2023using, guo2023identification, mishra2023comparative, mu2023densenet, rahman2023new]
Cerastium fontanum Common Mouse Ear Chickweed 1 [anderegg2023farm]
Chloris cucullata Hooded Windmill Grass 1 [saqib2023towards]
Cirsium arvense Creeping Thistle 5 [franco2018automatic, rasmussen2019pre, rasmussen2021pre, haq2023weed, saqib2023towards, thomas2023weakly]
Consolida regalis Forking Larkspur 3 [almalky2022efficient, almalky2023deep, almalky2023real]
Convolvulus arvensis Field Bindweed 3 [raptis2022multimodal, anderegg2023farm, saqib2023towards]
Dactylis glomerata Cat Grass 1 [saqib2023towards]
Echinochloa crus-galli Cockspur Grass 2 [xu2022multi, xu2023weedsnet]
Elymus repens Common Couch 1 [thomas2023weakly]
Equisetum arvense Horsetail 1 [zou2021modified]
Eschscholzia californica California poppy 1 [saqib2023towards]
Fallopia convolvulus Black Bindweed 1 [anderegg2023farm]
Fumaria officinalis Common fumitory 1 [anderegg2023farm]
Galium aparine Cleavers 34 [dyrmann2015roboweedsupport, dyrmann2017roboweedsupport, chavan2018agroavnet, dyrmann2018using, binguitcha2019crops, elnemr2019convolutional, jiang2019novel, espejo2020improving, fu2020field, rahman2020performance, skacev2020development, trong2020late, espejo2021combining, espejo2021testing, farkhani2021weed, mounashree2021automated, siddiqui2021neural, trong2021yielding, damilare2022weed, kumar2022weed, kundur2022deep, mu2022faster, reddy2022weed, srivastava2022role, tara2022machine, zhuang2022evaluation, anderegg2023farm, bharathi2023plant, dheeraj2023using, guo2023identification, liu2023semi, mishra2023comparative, mu2023densenet, rahman2023new]
Galium odoratum Sweet Woodruff 1 [liu2023semi]
Geranium pusillum Small-flowered Cranesbill 30 [dyrmann2017roboweedsupport, chavan2018agroavnet, dyrmann2018using, binguitcha2019crops, elnemr2019convolutional, jiang2019novel, espejo2020improving, fu2020field, rahman2020performance, skacev2020development, trong2020late, espejo2021combining, espejo2021testing, farkhani2021weed, mounashree2021automated, siddiqui2021neural, trong2021yielding, damilare2022weed, kumar2022weed, kundur2022deep, mu2022faster, reddy2022weed, srivastava2022role, tara2022machine, bharathi2023plant, dheeraj2023using, guo2023identification, mishra2023comparative, mu2023densenet, rahman2023new]
Humulus japonicus Japanese Hops 2 [wang2018low, wang2020transfer]
Lamium purpureum Red Deadnettle 2 [alirezazadeh2023comparative, anderegg2023farm]
Lolium multiflorum Italian Ryegrass 2 [jensen2021predictive, anderegg2023farm]
Lolium perenne Perennial Ryegrass 3 [dyrmann2017roboweedsupport, dyrmann2018using, saqib2023towards]
Lipandra polysperma Manyseed Goosefoot 1 [anderegg2023farm]
Lithospermum arvense Corn Gromwell 1 [alirezazadeh2023comparative]
Matricaria chamomilla Chamomile 1 [anderegg2023farm]
Matricaria discoidea Pineappleweed 2 [sukumar2016weed, pflanz2018weed]
Myosotis arvensis Field Forget-Me-Not 3 [dyrmann2017roboweedsupport, dyrmann2018using, anderegg2023farm]
Papaver rhoeas Common Poppy 3 [pflanz2018weed, anderegg2023farm, perez2023early]
Persicaria maculosa Redshank 3 [dyrmann2017roboweedsupport, dyrmann2018using, anderegg2023farm]
Phalaris paradoxa Awned Canarygrass 1 [jabir2021accuracy]
Poa annua Annual Bluegrass 3 [xu2020recognition, xu2022multi, anderegg2023farm]
Poa pratensis Kentucky bluegrass 1 [xu2023weedsnet]
Polygonum aviculare Common Knotweed 1 [anderegg2023farm]
Portulaca oleracea Purslane 1 [raptis2022multimodal]
Raphanus raphanistrum Wild Radish 2 [le2021detecting, shahbazi2021assessing]
Rumex crispus Curly Dock 2 [el2022metaheuristic, pp2023metaheuristic]
Rumex obtusifolius Broadleaf Dock 2 [el2022metaheuristic, pp2023metaheuristic]
Setaria adhaerens 1 [ronay2022effect]
Setaria viridis Green Foxtail 1 [zou2021modified]
Sinapis arvensis Charlock 31 [dyrmann2015roboweedsupport, dyrmann2017roboweedsupport, chavan2018agroavnet, dyrmann2018using, binguitcha2019crops, elnemr2019convolutional, jiang2019novel, espejo2020improving, fu2020field, rahman2020performance, skacev2020development, trong2020late, espejo2021combining, espejo2021testing, farkhani2021weed, mounashree2021automated, siddiqui2021neural, trong2021yielding, damilare2022weed, kumar2022weed, kundur2022deep, mu2022faster, reddy2022weed, srivastava2022role, tara2022machine, bharathi2023plant, dheeraj2023using, guo2023identification, mishra2023comparative, mu2023densenet, rahman2023new]
Sorghum halepense Johnson grass 1 [raptis2022multimodal]
Solanum nigrum Black Nightshade 4 [dyrmann2017roboweedsupport, dyrmann2018using, ronay2022effect, anderegg2023farm]
Spergula arvensis Corn Spurry 1 [alirezazadeh2023comparative]
Stellaria media Common Chickweed 35 [dyrmann2015roboweedsupport, dyrmann2017roboweedsupport, chavan2018agroavnet, dyrmann2018using, binguitcha2019crops, elnemr2019convolutional, jiang2019novel, espejo2020improving, fu2020field, rahman2020performance, skacev2020development, trong2020late, espejo2021combining, espejo2021testing, farkhani2021weed, mounashree2021automated, siddiqui2021neural, trong2021yielding, damilare2022weed, kumar2022weed, kundur2022deep, mu2022faster, reddy2022weed, srivastava2022role, tara2022machine, zhuang2022evaluation, alirezazadeh2023comparative, anderegg2023farm, bharathi2023plant, dheeraj2023using, guo2023identification, liu2023semi, mishra2023comparative, mu2023densenet, rahman2023new]
Taraxacum officinale Common Dandelion 1 [anderegg2023farm]
Trifolium pratense Red Clover 2 [dyrmann2017roboweedsupport, dyrmann2018using]
Tripleurospermum inodorum Scentless Mayweed 32 [dyrmann2015roboweedsupport, dyrmann2017roboweedsupport, chavan2018agroavnet, dyrmann2018using, binguitcha2019crops, elnemr2019convolutional, jiang2019novel, espejo2020improving, fu2020field, rahman2020performance, skacev2020development, trong2020late, espejo2021combining, espejo2021testing, farkhani2021weed, mounashree2021automated, siddiqui2021neural, trong2021yielding, damilare2022weed, kumar2022weed, kundur2022deep, mu2022faster, reddy2022weed, srivastava2022role, tara2022machine, bharathi2023plant, dheeraj2023using, guo2023identification, mishra2023comparative, mu2023densenet, rahman2023new, thomas2023weakly]
Urtica urens Small Nettle 2 [dyrmann2017roboweedsupport, dyrmann2018using]
Veronica hederifolia Ivy-leaved Speedwell 2 [pflanz2018weed, anderegg2023farm]
Veronica persica Common field-speedwell 5 [sukumar2016weed, dyrmann2017roboweedsupport, dyrmann2018using, anderegg2023farm, liu2023semi]
Viola arvensis Field Pansy 4 [dyrmann2017roboweedsupport, dyrmann2018using, pflanz2018weed, anderegg2023farm]
Vulpia myuros Rattail Fescue 1 [jensen2021predictive]
Xanthium strumarium 2 [wang2018low, wang2020transfer]
[…]